OpenStudy (anonymous):

Two similar right triangles have areas of 6 square inches and

Directrix (directrix):

and ...?

OpenStudy (anonymous):

150 square inches. The length of the hypotenuse of the smaller triangle is 5 inches. What is the sum of the lengths of the legs of the larger triangle?

Directrix (directrix):

The ratio of the corresponding area from small to large = 6/150 = 1/25. For similar triangles, the square of the scale factor is the ratio of corresponding linear parts of the 2 trianlges. So the scale factor is 1 to 5 or 1/5. and....

Directrix (directrix):

1/5 = 5/h where h is the hypotenuse of the larger triangle. h = 25.

Directrix (directrix):

The perimeters are also in the ratio 1/5 -- same as scale factor.

Directrix (directrix):

What is the sum of the lengths of the legs of the larger triangle? If the sides of the larger triangle form a Pythagorean triple, then that triple is (9, 12, 15) which is a multiple of (3,45). 9+12+15 = 36 - 15 = 21 for sum of legs.